# Cycle Consistency as Reward: Learning Image-Text Alignment without Human Preferences

### v1

@@ 3.3 Reward Modeling
GPT-4o 61.3% 60.0% 48.1% 45.8% 24.8%
Raw Cycle Consistency 58.6% 61.2% 60.5% 59.8% 54.5%
CycleReward-I2T 63.9% 65.6% 66.5% 65.7% 60.20%
CycleReward-T2I 57.1% 78.2% 68.3% 66.2% 60.2 %
CycleReward-Combo 66.5% 63.8% 67.7% 65.8% 61.3%
Table 2: Agreement rates between human preferences and those derived from GPT-4o, raw cycle consistency, and CycleReward.
4 Does cycle consistency align with human preferences?
Table 2 directly compares annotating preferences with cycle consistency to human labels. We measure the agreement rate between cycle consistency and human preferences on detailed captioning and text-to-image generation. For detailed captioning, we compare to preferences from the RLHF-V [96] and POVID [105] datasets. For text-to-image generation, we compare to preferences from Human Preferences Dataset v2 (HPDv2) [88], Pick-a-Pic v2 [40], and ImageRewardDB [91]. For each dataset, we sample 1K random binary comparison pairs. We compare preferences from both “raw cycle consistency” scores $s(x\rightarrow y)$, $s(y\rightarrow x)$, and our trained reward models. We also include a comparison with GPT-4o [61] annotations as they have been shown to be useful for preference learning [48].
CycleReward achieves the highest agreement with human annotations, with CycleReward-Combo having the highest average agreement rate of 65%. While GPT-4o annotations align more closely with human preferences on detailed captioning tasks, its agreement drops significantly on text-to-image generation, with as low as 24.84% on ImageRewardDB. In contrast, raw cycle consistency maintains a consistent agreement rate across both tasks. Training a reward model on cycle consistency further improves alignment, which demonstrates the effectiveness of distilling cycle-consistent preferences into a learned reward model.
While we compare with human preferences, our aim is not to mimic them. We aim to learn image-text alignment, and demonstrate that cycle consistency is an effective proxy — achieving strong results without collecting any human labels, as shown in the following sections.
@@ 4 Does cycle consistency align with human preferences?
Table 3: Evaluating image-text alignment. CycleReward-Combo and CycleReward-I2T outperform all approaches on detailed captioning evaluation, even those trained on human preferences. Notably, we outperform VQAScore x24 larger in model size. For text-to-image generation, CycleReward achieves similar performance to models trained on human preferences, while VQAScore outperforms others. Across both tasks, our learned reward model outperforms using raw cycle consistency.
5 Reward Model Evaluation
We evaluate CycleReward’s ability to assess and improve image-text alignment across two tasks: detailed captioning and text-to-image generation. Specifically, we evaluate CycleReward as an alignment metric, and then deploy it to maximize inference time alignment via Best-of-$N$.
Comparison methods.
We compare against existing reference-free metrics for image-text alignment. These include: (1) CLIPScore [29] which measures cosine similarity between image and text embeddings from CLIP [69], (2) ImageReward [91], (3) HPSv2 [89, 88] and (4) PickScore [40], which are trained on large human preference datasets for text-to-image generation, and (5) VQAScore [52] which produces alignment scores by querying a VLM with the prompt "Does this figure show {text}?". For VQAScore, we compare two different model sizes: CLIP-T5-xl (3B) and CLIP-T5-xxl (11B). (6) Raw cycle consistency directly measures alignment scores $s(x\rightarrow y)$ for image-to-text generation and $s(y\rightarrow x)$ for text-to-image generation without learning a reward model. For image-to-text generation, this is equivalent to Image2Text2Image [32]. In our experiments, we adopt the same model configurations (i.e., choice of decoders, similarity metrics) for a fair comparison.
Figure 3: Best-of-$N$ relative performance gain. From left to right: LLaVA-W, DeCapBench, T2I-CompBench (mean of 6 categories), and PartiPrompts (complex). In each plot, we show the relative performance gain from BoN sampling with different metrics. Feedback from our reward model leads to the greatest overall improvement for detail captioning tasks, while we maintain competitive text-to-image generation performance with VQA-Score and ImageReward.
@@ 5.1 Metric for Image-Text Alignment
While many benchmarks exist for short captions, few target detailed image captioning, and those that do often lack labels or contain limited examples. One exception is DetailCaps-4870 [16], which evaluates captions on accuracy and inclusion across object, attribute, and relation categories. It contains 4,870 image-text pairs from ShareGPT4V [8], LLaVA 1.5, and CogVLM [31, 87], scored by three VLMs: GPT-4V [62], Gemini-1.5 Pro [82], and GPT-4o [61]. We use the mean score as a pseudo-ground truth, and measure agreement with alignment metrics using pairwise accuracy [13]. To evaluate text-to-image generation, we use GenAI-Bench [52, 43], which consists of 1,600 prompts paired with 6 generated images from different models. Each generation is annotated with three human ratings based fidelity to the text.
@@
We directly compare to ImageReward, which uses human labels, by training a reward model on the same backbone and ImageRewardDB dataset but re-annotated with cycle consistency preferences. We refer to this model as IRDB-Cycle. Table 3 shows IRDB-Cycle achieves comparable performance to ImageReward, demonstrating that cycle consistency is an effective and cheaply scalable alternative for human labels. In the following sections, we show training on our dataset yields further improvements.
Evaluating detailed captioning.
Table 3 shows that CycleReward outperforms all existing methods by a large margin, including HPSv2, PickScore, and ImageReward, which are trained on human preferences. Notably, CycleReward outperforms VQAScore-11B by 10.26%, which is 24x larger in model size. It outperforms raw cycle consistency, i.e., computing cycle consistency on-the-fly as in Image2Text2Image [32], which highlights the effectiveness of distilling cycle consistency into a learned reward model.
Evaluating text-to-image generation.
Table 3 reports pairwise accuracy between different methods and human preferences on GenAI-Bench. CycleReward performs comparably to HPS, PickScore, and Imagereward — models trained with human annotations. CycleReward outperforms both raw cycle consistency and IRDB-Cycle, a model trained on ImageRewardDB with cycle-consistent labels. Although VQAScore (11B) aligns most with humans, our model does surprisingly well considering its small scale (477M). See Appendix E.1 for qualitative comparisons.
@@ 5.2 Best-of-N Sampling
Best-of-$N$ (BoN) sampling is a simple strategy to improve models at test time [10, 79, 58]. The process involves generating $N$ candidate outputs from a base model which are then ranked according to a reward model which selects the one with the highest score. The selection criterion is entirely based on the reward model, and naturally better models choose higher-quality outputs.
@@
For image-to-text generation, we use two detailed captioning benchmarks: LLaVA-W [53] detailed captioning subset and DeCapBench [92], which assess the correctness and coverage of details (i.e., precision and recall) in generated captions. LLaVA-WD evaluations are conducted using GPT-4o-mini [61] as the evaluator model, while DeCapBench uses DCScore [92]. For text-to-image generation, we use T2I-Compbench [33] for fine-grained preferences on six compositional categories, and the “Complex” subset of PartiPrompts [95] for complex, detailed prompts.
Improving detailed captioning.
For each image, we perform BoN selection from a pool of 250 captions obtained from a combination of temperature, nucleus, and prompt sampling LLaVA1.5-13B [44, 55]. For image captioning tasks, BoN sampling with our reward model increases performance significantly over other metrics as seen in 3. Both LLaVA-W and DeCapBench assess captions based on correctness and level of detail, and our reward model yields the largest improvement in the overall evaluation score. For further analysis we plot BoN results for the non-hallucination and comprehensiveness scores from DeCapBench and find that our model excels at describing many things in detail while maintaining correctness (albeit less accurately than VQAScore). In contrast, baselines such as VQAScore and ImageReward highly weigh accuracy to the point of preferring captions with significantly less detail.
Improving text-to-image generation. For all text prompts we use SDXL-Turbo [74] to generate a pool of 100 images with different random seeds to perform BoN sampling. Figure 3 (right) shows relative performance gain using different reward models for BoN sampling. Note that our self-supervised reward models perform similarly to ImageReward which is trained with human preferences, and even outperforms on ‘complex’ text prompts. For specific categories in T2I-CompBench see E.2.
@@ 5.3 Ablation Study
We ablate several design choices for our reward model on DetailCaps-4870 and GenAI-Bench. For both benchmarks, we report pairwise accuracy, and gray entries denote design choices used in CycleReward.
@@
Table 5: Effect of decoders for generating reconstructions. Choices used by our model are in gray.
Ablation DetailCaps-4870 GenAI-Bench
Best variant (CR-I2T) 58.02 53.49
MSE loss 41.87 40.57
1K images 52.86 44.39
Without filtering 57.28 51.92
@@
Table 7: Direct preference optimization (DPO) for image-to-text generation. The best results are indicated in bold. DPO with CycleReward-VL improves the base model’s performance across all tasks – including detailed captioning, perception, reasoning, and hallucination reduction – despite only containing captioning instructions. It achieves comparable or superior results to VLFeedback, a preference dataset annotated with GPT-4V spanning diverse task instructions.
@@
Model Spatial Color Complex Numeracy Shape Texture DrawBench PP-Simple Detail PP-FG Detail PP-Complex
@@
Table 8: Direct preference optimization (DPO) for text-to-image generation. For all evaluations, higher scores are better. T2I-Compbench and DrawBench scores range from 0 to 100 while PartiPrompt (PP) scores range from 1 to 10. In all cases, the Diffusion DPO training with CycleReward-VL outperforms the base model. Furthermore, our model often outperforms or is comparable with the Pick-A-Pic Diffusion DPO model, especially for longer text prompts.
Objective function, data scale, and filtering.
@@ 6.1 Results
To our surprise, DPO finetuning with CyclePrefDB-I2T enhances the base model’s performance across all vision-language tasks – including detailed captioning, perception, reasoning, and hallucination – although our dataset only contains captioning instructions. Despite our narrow task instruction and smaller size, it achieves comparable or superior results to VLFeedback, a preference dataset annotated by GPT-4V across VQA, captioning, classification, reasoning, conversation, and red teaming instructions.
@@
Table 8 reports evaluation results on T2I-CompBench and DrawBench (scores from 1 to 100) and PartiPrompts (scores from 1 to 10), where higher is better. Across all categories, the model trained on CyclePrefDB-T2I outperforms the base SD1.5 model and is comparable with or outperforms the Pick-A-Pic model on complex prompts, which is particularly a challenge for SD1.5. Qualitative results are shown in the Appendix 14.
@@ 7 Discussion
Our findings demonstrate that cycle consistency provides a scalable and effective supervisory signal for image-text alignment, achieving competitive performance without relying on any human-labeled data. By constructing CyclePrefDB, a preference dataset annotated via cycle consistency, we enable training reward models that generalize across both image-to-text and text-to-image tasks. These models outperform or match existing baselines on detailed captioning and compositional text-to-image benchmarks, suggesting that cycle consistency can serve as an effective alternative to human or AI-based preference annotations.
Figure 5: Failure cases. (Left): Despite being faithful to the input text, the right image is rejected as the reconstructed text contains hallucination inconsistent with the original prompt. (Right): The short caption is preferred over the descriptive caption due to an error in text-to-image generation. We show the reconstructed image under each caption.
However, cycle consistency is not without limitations. The quality of supervision depends on accurate reconstructions from pre-trained decoders, and errors in generation can lead to misleading preference signals (Figure 5). Our method also inherits biases from the underlying models used for reconstructions and and similarity measurements. Stable Diffusion 3 has a 77-token limit which limits consideration of longer texts, and LLaVA-1.5-3B can be prone to hallucinations. DreamSim often favors images with similar foregrounds over backgrounds [22], and SBERT is sensitive to text style. Furthermore, we observe worse text-to-image performance, which may partially stem from dataset differences. HPSv2, PickScore, and ImageReward are trained on prompts from real users often describing artwork, whereas CycleReward is trained on LLM-summarized descriptions for natural images. Moreover, cycle consistency primarily considers preservation of information, while other aspects such as aesthetics or style may also affect human preferences. Future work could address these challenges by improving reconstruction quality, prompt diversity, and applying cycle consistency in different scenarios. More broadly, our framework offers a general approach for learning dense alignment between two modalities, and could be extended to other domains such as audio-text, video-language, or even reasoning tasks.
@@ Appendix A Cycle Consistency and Point-wise Mutual Information
In Equation 3, we define $s(x\rightarrow y)$ and $s(y\rightarrow x)$ with respect to fixed backward mappings $G:Y\rightarrow X$ and $F:X\rightarrow Y$ respectively. If $F,G$ are stochastic mappings, then we can view $G$ as sampling some $x^{\prime}=G(y)$ from the distribution $p_{G}(X|Y=y)$ a distribution which is determined by $G$. Symmetrically, we can view $F$ as sampling $y^{\prime}=F(x)$ from the distribution $p_{F}(Y|X=x)$ determined by $F$. We then argue that distributionally,
@@
$\begin{split}s&(x,y)_{d}=\log p(x|y)+\log p(y|x)\\ &=\log\frac{p(x,y)}{p(y)}+\log\frac{p(x,y)}{p(x)}\\ &=\log\frac{p(x,y)^{2}}{p(x)p(y)}\\ &=\log p(x,y)+\text{PMI}(x,y)\\ \end{split}$ (10)
@@ Appendix B Benefits from Reward Modeling
Figure 6 plots DetailCaps-4870 benchmark performance against the number of samples $N$ used for mean cycle consistency score. Although using more seeds generally benefits raw cycle consistency, improvement tapers off around $N=5$ and never reaches the performance of CycleReward.
@@ Appendix C CyclePrefDB Dataset Details
We provide examples of reconstructed images and texts used to create comparison pairs in our dataset in Figure 7. Generally, we find that better, more descriptive image captions lead to image reconstructions that are more similar to the input image. Symmetrically, generated images that are faithful to the prompt have text reconstructions that reflect this. However, there exists failure cases often due to poor reconstructions as in Figure 5.
@@
Common strategies for filtering human preferences include: (1) removing duplicate entries, (2) filtering out cases where both responses are harmful or irrelevant [93], and (3) excluding low-margin examples where one response is only marginally better than the other [12]. Following these principles, we adopt a similar filtering strategy by removing duplicate captions, excluding examples where the reward difference is within a certain threshold, i.e., $|r_{i}-r_{j}|<\tau_{\text{sim}}$, and discarding comparison pairs where the preferred reward is below a threshold, i.e., $r_{i}<\tau_{\text{neg}}$. In practice, we use $\tau_{\text{sim}}=0.005$, $\tau_{\text{neg}}=0.7$ for DreamSim, and $\tau_{\text{neg}}=0.4$ for SBERT. In practice, training with this dataset filtering leads to a small performance gain on alignment benchmarks and a bigger performance gap in best-of-$N$ experiments seen in Figure 9.
Figure 9: Best-of-$N$ results with and without dataset filtering. Filtering the dataset improves inference-time optimization by enabling better candidate selection during best-of-$N$ sampling.

### v2

@@ 4 Cycle Consistency and Human Preferences
GPT-4o 61.3 60.0 48.1 45.8 24.8
Raw Cycle Consistency 58.6 61.2 60.5 59.8 54.5
CycleReward-I2T 63.9 65.6 66.5 65.7 60.2
CycleReward-T2I 57.1 78.2 68.3 66.2 60.2
CycleReward-Combo 66.5 63.8 67.7 65.8 61.3
Table 2: Agreement rates (%) between human preferences and those from GPT-4o, raw cycle consistency, and CycleReward.
5 Reward Model Evaluation
We evaluate CycleReward’s ability to assess and improve image-text alignment across two tasks: detailed captioning and text-to-image generation. Specifically, we evaluate CycleReward as an alignment metric, and then deploy it to maximize inference-time alignment via Best-of-$N$.
Comparison methods.
We compare against current reference-free image-text alignment metrics. These include: (1) CLIPScore [29] which measures cosine similarity between image and text embeddings from CLIP [68], (2) ImageReward [91], (3) HPSv2 [89, 88] and (4) PickScore [40], which are trained on large human preference datasets for text-to-image generation, and (5) VQAScore [51] which produces alignment scores by querying a VLM with the prompt "Does this figure show {text}?". For VQAScore, we compare two different model sizes: CLIP-T5-xl (3B) and CLIP-T5-xxl (11B). (6) Raw cycle consistency directly uses alignment scores $s(x\rightarrow y)$ for image-to-text generation and $s(y\rightarrow x)$ for text-to-image generation without learning a reward model. For image-to-text, this is equivalent to Image2Text2Image [32]. We adopt the same model configurations (i.e., decoders, similarity metrics) for fair comparison.
@@ 5 Reward Model Evaluation
[…]
Figure 3: Best-of-NN relative performance gain. From left to right: LLaVA-W, DeCapBench, T2I-CompBench (mean of 6 categories), and PartiPrompts (complex). In each plot, we show the relative performance gain from BoN sampling with different metrics. Feedback from our reward model leads to the greatest overall improvement for detailed captioning tasks, while we maintain competitive text-to-image generation performance with VQAScore and ImageReward.
@@ 5.1 Metric for Image-Text Alignment
While many benchmarks exist for short captions, few target detailed descriptions, and those that do often lack labels or contain limited examples. One exception is DetailCaps-4870 [16], which evaluates captions on accuracy and inclusion across object, attribute, and relation categories. It contains 4,870 image-text pairs from ShareGPT4V [8], LLaVA 1.5, and CogVLM [31, 87], scored by three VLMs: GPT-4V [61], Gemini-1.5 Pro [82], and GPT-4o [60]. We use the mean score as a pseudo-ground truth. To evaluate text-to-image generation, we use GenAI-Bench [51, 43], which consists of 1,600 prompts paired with 6 generated images from different models. Each generation is annotated with three human ratings based on fidelity to the text. For both tasks, we measure agreement with alignment metrics using pairwise accuracy [13].
@@
We directly compare to ImageReward, which uses human labels, by training a reward model on the same backbone and ImageRewardDB dataset, but re-annotated with cycle consistency preferences. We refer to this model as IRDB-Cycle. Table 3 shows IRDB-Cycle achieves comparable performance to ImageReward, demonstrating that cycle consistency is an effective and cheaply scalable alternative for human labels. In the following sections, we show training on our dataset yields further improvements.
Results.
[…]
For text-to-image generation, CycleReward performs comparably to HPS, PickScore, and ImageReward, all of which are trained with human annotations. CycleReward outperforms both raw cycle consistency and IRDB-Cycle, a model trained on ImageRewardDB with cycle-consistent labels. Although VQAScore (11B) aligns most with humans, our model does surprisingly well considering its small scale (477M). See Appendix E.1 for qualitative comparisons.
@@ 5.2 Best-of-N Sampling
Best-of-NN (BoN) sampling is a simple strategy to improve model results at test time [10, 79, 57]. The process involves generating $N$ candidate outputs from a base model, ranking them using a reward model, and selecting the one with the highest score. The selection criterion is entirely based on the reward model, and naturally better models choose higher-quality outputs.
@@
For image-to-text generation, we use two detailed captioning benchmarks: LLaVA-W [52] detailed captioning subset (LLaVA-WD) and DeCapBench [92], which assess the correctness and coverage of details (i.e., precision and recall) in generated captions. LLaVA-WD evaluations are conducted using GPT-4o-mini [60] as the evaluator model, while DeCapBench uses DCScore [92]. For text-to-image generation, we use T2I-Compbench [33] for fine-grained preferences on six compositional categories, and the “complex” subset of PartiPrompts [95] for complex, detailed prompts.
Detailed captioning results.
For each image, we perform BoN selection from a pool of 250 captions obtained from a combination of temperature, nucleus, and prompt sampling LLaVA1.5-13B [44, 54] (see Appendix E.2 for details). For image captioning, BoN sampling with our reward model increases performance significantly over other metrics as seen in Figure 3. Both LLaVA-WD and DeCapBench assess captions based on correctness and level of detail, and our reward model yields the largest improvement in the overall evaluation score. In Appendix E.2 we plot BoN results for the non-hallucination and comprehensiveness scores from DeCapBench and find that our model excels at describing many things in detail while maintaining correctness (albeit less accurately than VQAScore). In contrast, baselines such as VQAScore and ImageReward highly weigh accuracy to the point of preferring captions with significantly less detail.
Text-to-image generation results. For all text prompts, we use SDXL-Turbo [74] to generate a pool of 100 images with different random seeds to perform BoN sampling. Figure 3 (right) shows relative performance gain using different reward models for BoN sampling. Note that our self-supervised reward models perform similarly to ImageReward which is trained with human preferences, and even outperforms on “complex” text prompts. For specific T2I-CompBench category results see Appendix E.2.
@@ 5.3 Ablation Study
We ablate several design choices for our reward model on DetailCaps-4870 and GenAI-Bench. For both benchmarks, we report pairwise accuracy, and gray entries denote design choices used in CycleReward. For ablations on objective function, data scale, and filtering see Appendix E.4.
@@
Table 5: Effect of decoder models for generating reconstructions. Choices used by our model are in gray.
@@
Table 6: Direct preference optimization (DPO) for image-to-text generation. The best results are indicated in bold. DPO with CyclePrefDB-I2T improves the base model’s performance across all tasks—including detailed captioning, perception, reasoning, and hallucination reduction—despite only containing captioning instructions. It achieves comparable or superior results to VLFeedback, a preference dataset annotated with GPT-4V spanning diverse task instructions.
@@
Model Spatial Color Complex Numeracy Shape Texture DrawBench PP-Simple Detail PP-FG Detail $P$-Complex
@@
Table 7: Direct preference optimization (DPO) for text-to-image generation. For all evaluations, higher scores are better. T2I-Compbench and DrawBench scores range from 0 to 100 while PartiPrompt ($P$) scores range from 1 to 10. In all cases, the Diffusion DPO training with CyclePrefDB-T2I outperforms the base model. Furthermore, our model often outperforms or is comparable with the Pick-A-Pic Diffusion DPO model, especially for longer text prompts.
@@ 6.1 Results
To our surprise, DPO fine-tuning with CyclePrefDB-I2T enhances the base model’s performance across all vision-language tasks—including detailed captioning, perception, reasoning, and hallucination—although our dataset only contains captioning instructions. Despite our narrow task instruction and smaller dataset size, it achieves comparable or superior results to VLFeedback, a preference dataset annotated by GPT-4V across VQA, captioning, classification, reasoning, conversation, and red teaming instructions.
@@
Table 7 reports evaluation results on T2I-CompBench and DrawBench (scores from 1 to 100) and PartiPrompts (scores from 1 to 10), where higher is better. Across all categories, the model trained on CyclePrefDB-T2I outperforms the base model and is comparable with or outperforms the Pick-A-Pic model especially on complex prompts, which is particularly a challenge for Stable Diffusion 1.5. See Figure 4 and Appendix 14 for qualitative results.
@@ 7 Discussion
We find that cycle consistency provides a scalable and effective supervisory signal for image-text alignment, achieving competitive performance without relying on any human-labeled data. We first construct CyclePrefDB, a preference dataset annotated via cycle consistency, and then train reward models that generalize across both image-to-text and text-to-image tasks. These models outperform or match existing baselines on detailed captioning and compositional text-to-image benchmarks, suggesting that cycle consistency is an effective alternative to human annotations.
However, our method has limitations. Supervision quality depends on accurate reconstructions from pre-trained decoders, and generation errors can mislead preferences. Appendix E.6 visualizes failures cases and discusses more limitations. Future work could address these challenges by improving reconstructions, prompt diversity, and applying cycle consistency in different scenarios. Broadly, our framework offers a general approach for learning dense alignment between modalities, and could be extended to new domains such as audio-text, video-language, or even reasoning tasks.
@@ Appendix A Cycle Consistency and Point-wise Mutual Information
In Equation 3, we define $s(x\rightarrow y)$ and $s(y\rightarrow x)$ with respect to fixed backward mappings $G:Y\rightarrow X$ and $F:X\rightarrow Y$ respectively. If $F,G$ are stochastic mappings, then we can view $G$ as sampling some $x^{\prime}=G(y)$ from the distribution $p_{G}(X|Y=y)$ - a distribution which is determined by $G$. Symmetrically, we can view $F$ as sampling $y^{\prime}=F(x)$ from the distribution $p_{F}(Y|X=x)$ determined by $F$. We then argue that distributionally,
@@
[…]
@@ Appendix B Benefits from Reward Modeling
Figure 5 plots DetailCaps-4870 benchmark performance against the number of samples $N$ used to compute the mean cycle consistency score. Although using more seeds benefits raw cycle consistency, improvement tapers off around $N=5$ and never reaches the performance of CycleReward.
@@ Appendix C CyclePrefDB Dataset Details
We provide examples of reconstructed images and texts used to create comparison pairs in our dataset in Figure 6. Generally, we find that better, more descriptive image captions lead to image reconstructions that are more similar to the input image. Symmetrically, generated images that are faithful to the prompt have text reconstructions reflecting this. However, failure cases can occur due to poor reconstructions as in Figure 15.
@@
Common strategies for filtering human preferences include: (1) removing duplicate entries, (2) filtering out cases where both responses are harmful or irrelevant [93], and (3) excluding low-margin examples where one response is only marginally better than the other [12]. Following these principles, we adopt a similar filtering strategy by removing duplicate captions, excluding examples where the reward difference is within a certain threshold, i.e., $|r_{i}-r_{j}|<\tau_{\text{sim}}$, and discarding comparison pairs where the preferred reward is below a threshold, i.e., $r_{i}<\tau_{\text{neg}}$. In practice, we use $\tau_{\text{sim}}=0.005$, $\tau_{\text{neg}}=0.7$ for DreamSim, and $\tau_{\text{neg}}=0.4$ for SBERT. In practice, training with dataset filtering leads to a small performance gain on alignment benchmarks and a bigger performance gap in Best-of-$N$ experiments as seen in Figure 8.
Figure 8: Best-of-NN results with and without dataset filtering. Filtering the dataset improves inference-time optimization by enabling better candidate selection during best-of-NN sampling.